OpenStudy (anonymous):

find the value of x and y. top= 3x + 17 bottom= 2y left side= 4 right side= 4x - y

OpenStudy (fj):

x=5 y=16

OpenStudy (anonymous):

thanks but how did you get that

OpenStudy (fj):

you have equale top whit bottom and left side whit right side too.

OpenStudy (anonymous):

so 3x + 17= 2y

OpenStudy (fj):

that's right

OpenStudy (anonymous):

how do i equal out the x and y

OpenStudy (fj):

you have to solve this: 4=4x-y and 2y=3x+17 ok?

OpenStudy (fj):

\[3x-2y =-17\] \[-2\times \left( 4x-y=4 \right)\] \[-8x+2y=-8\] \[3x-2y=-17\] \[x=5 , y=16\]

OpenStudy (fj):

you got it?

OpenStudy (anonymous):

yeah kinda thanks